# EgoOops: A Dataset for Mistake Action Detection from Egocentric Videos referring to Procedural Texts

### v2

@@ Abstract
Fig. 1: An example task (cardboard) from our EgoOops dataset. EgoOops includes 50 egocentric videos across five procedural domains and corresponding procedural texts. It contains three types of annotations: video-text alignment, mistake labels, and descriptions explaining the errors in each segment.
Index Terms—  Mistake detection, Procedural activity, Video dataset, Egocentric video
@@ 1 Introduction
Table 1: Comparison of mistake datasets. Range: mistake labels to a video or each step segment with start and end times; Cat.: finer-grained categorization than correct , mistake , and correction ; Desc.: descriptions explaining why each segment is incorrect; OM, EM: order mistakes, execution mistakes (see Section 3.1); Proc.: workers follow step-by-step procedural texts. Domain specific (e.g., extra screws). Orders unseen in the training set, not always faulty.
@@
However, these studies have the following three limitations (see Table 1). L1: video-focused approaches.Existing studies [1, 2, 6, 12, 13, 5] have focused on video-only approaches to mistake detection and not utilized procedural texts. Their approaches are suitable for visually apparent mistakes like incorrectly attached parts and falling plates. Nevertheless, in text-following activities, some mistakes are deviations from procedural texts, thus not obvious only from visual cues (e.g., in Fig. 1, the use of tape is a mistake because the text designates glue). Therefore, besides videos, texts are essential for models to detect mistakes accurately. L2: specific domains.Procedural texts are included in only a few mistake datasets recording cooking [6, 5]. Since many real-world activities follow procedural texts, it is essential to collect data from more domains. L3: rough mistake labels.Most datasets define coarse-grained (correct /mistake /correction[1, 3, 4] or domain-specific (e.g., extra screws[2, 6] categories. General and fine-grained categorization enables analysis of mistake patterns across diverse domains (e.g., a commonly frequent category of mistakes).
@@
We also propose an approach to the problem of mistake action detection, especially focusing on the utilization of procedural texts (L1). To leverage the texts, our approach combines video-text alignment and mistake label classification; the former localizes the start and end times of each procedural step, and the latter assigns the step segments one label of mistake classes. In experiments using EgoOops, our multi-modal approach outperforms a video-only baseline, and the ablation of textual inputs decreases our performance. These results demonstrate that incorporating procedural texts is essential for mistake detection.
@@ 2 Related work
Mistakes are crucial issues in procedural activities, leading to recent proposals of various datasets (see Table 1). These are categorized into free-style and text-following settings.
@@ 3 EgoOops dataset
Our dataset consists of procedural texts and egocentric videos containing mistake actions. This section describes terms, text preparation, recording process, annotations, and statistics.
@@ 3.1 Mistake definition
We define mistakes as deviations from a procedural text. This definition leads to two mistake types: order mistakes and execution mistakes . Order mistakes occur when the steps executed by a worker are different from the ones listed in the procedural text. These include skipping necessary steps (missing ), swapping the step order (out-of-order ), inserting extra steps (undefined ), and pausing a step and resuming after others (split ). Execution mistakes occur when a worker fails to follow the instructions while carrying out a step. We categorize them into the following six classes: 1. working with wrong objects (object ), 2. grasping wrong objects and releasing them without using (mispick ), 3. correction of mistake actions (correction ), 4. unintended actions (accident ), 5. performing in the wrong way (way ), 6. others (others ).
3.2 Task selection and text preparation
To record procedural activities in diverse domains with various mistake actions, we selected the following five tasks: electrical circuits (EC ), color mixture experiments (CM ), ionic reaction experiments (IR ), toy building blocks (BB ), and cardboard crafts (CB ). Their domains are diverse (i.e., electronics, chemistry, assembly, and crafts), and various mistakes can happen (e.g., using wrong chemicals, accidentally cutting off cardboard).
One of the authors wrote a procedural text for each task before recording. They referred to the web for color mixture and cardboard crafts, and texts for ionic reactions and circuits are borrowed from out-of-the-box kits. For building blocks, they wrote it from scratch because we found no references. Note that all texts were originally written in Japanese, then we manually translated them to prepare an English version.
@@ 3.3 Video recording
Four Japanese graduate students (4 males) performed the tasks following the texts while equipped with a head-mounted camera. A participant worked on every task two or four times, totaling 10 recordings for each task. To record various mistake actions, they intentionally performed mistake actions that we assumed in advance (e.g., skipping several steps). Out of ten times to record videos, they contained the intentional mistakes five times and tried to follow procedures correctly for the other five times. Note that accidental mistakes due to carelessness also happened even in the latter.
@@ 3.4 Annotations
EgoOops contains annotations of video-text alignment, mistake labels, and their descriptions. One person annotates the whole dataset using a web annotation tool described in the supplementary materials.
We first annotate video-text alignment by extracting start and end timestamps (i.e., segments) and mapping them to the corresponding steps in a procedural text. Note that we also annotate steps not written in the procedural text as undefined (e.g., grasping the wrong object). The video-text alignment itself reveals order mistakes. For example, if step 2 occurs before step 1, these steps are considered swapped. Thus, we do not explicitly annotate order mistakes.
If a segment indicates an execution mistake, we assign one label of the six execution mistake classes listed in Section 3.1. This categorization can be used to analyze mistake patterns across diverse tasks (e.g., a commonly frequent category of mistakes). In addition, we also provide the mistake segments with descriptions of why the performed actions are considered mistakes. They can be used to analyze model performance in detail or interact with workers through smart assistants.
We ensure the quality of the original annotations by testing for agreement with another annotator. For video-text alignment, the annotator newly extracts segments and maps them to the corresponding step labels, resulting in the temporal Intersection over the Union (tIoU) to the original at 88.8. In addition, they watch the original segments and give new mistake labels and descriptions. The labels and descriptions are compared using Cohen’s kappa [14] and BERTScore [15], achieving 86.8 and 96.3, respectively. These high scores confirm the quality of the original annotations.
@@ 3.5 Statistics
Table 2 shows different trends between the tasks in terms of video duration, segment duration, and the number of segments. For video duration, the longest is the cardboard task at 26.1 minutes, while the shortest is the building block task at 1.9 minutes. As for the texts, Table 2 compares the number of steps in a procedural text and the number of words per step. The task with the most steps is the cardboard task, while the task with the longest instructions is the building block task. These statistics indicate that EgoOops covers a variety of procedural tasks ranging from short to long.
@@
#vid. Avg. (min) #seg. #seg. / #vid. Avg. (sec) #steps per text #words / #steps
@@
Table 3 shows comparisons of execution and order mistakes between tasks. Of all the mistakes, 41% ($=95/233$) are execution mistakes, and the others are order mistakes. Among the order mistakes, splitting steps is the most frequent. Table 4 shows the counts of the labels for execution mistakes. The two most frequent labels are grasping the wrong objects and releasing them without use (label 2) and working in the wrong way or moving (label 5). In addition, we also find unique mistake patterns of each task (e.g., unintended actions (label 4) frequently happen in ionic reaction experiments but rarely in the other tasks).
Table 3: Comparisons of execution and order mistakes.
Task Execution mistake Order mistake
Missing Out-of-order Split Undefined
EC 22 2 10 12 6
CM 22 1 2 3 9
IR 19 2 4 1 6
BB 19 0 3 6 10
CB 13 7 21 29 4
Total 95 12 40 51 35
Table 4: Statistics of mistake labels.
@@ 4 Application: mistake action detection
We propose a text-oriented approach to mistake action detection. This section formalizes the problem, explains our approach, and reports experimental results on EgoOops.
@@ 4.1 Problem formalization
Mistake action detection consists of two problems: video-text alignment and mistake label classification. It finally localizes temporal segments and classes of mistakes in a video.
@@
The alignment outputs are passed to the next problem of mistake label classification, which predicts a mistake label for each step segment. The $k$-th step segment is represented by extracted video frames based on the start and end frame numbers as $\mathbf{V}^{\prime}_{k}=(\mathbf{f}_{s_{k}},\dots,\mathbf{f}_{s_{e}})$. Our objective is to assign the segment $\mathbf{V}^{\prime}_{k}$ one label of mistake classes. In our experiments, we group the mistakes except for 3. correction (see Section 3.1) into the common mistake class and solve the classification of three classes: correct , mistake , and correction .  [^1] Our preliminary experiments showed that distinguishing the seven classes (six classes in Section 3.1 plus the correct class) is difficult currently.
4.2 Models
Video-text alignment. We adopt StepFormer [16] to tackle video-text alignment. Stepformer was originally proposed to learn video-text alignment from untrimmed videos with narrations in a self-supervised fashion [16]. We train it to fit EgoOops better in the following two steps. First, we apply its self-supervised learning to pre-train it on Ego4D [10] using EgoVLPv2 [17] features. Second, we add a fully-supervised loss to learn from video-text alignment annotations and fine-tune the pre-trained model on EgoOops. Specifically, we improve the similarity between the remaining slots [16] and annotated video segments, and vice versa, using InfoNCE contrastive loss [18]. We call the resulting model StepFormer++ to distinguish it from the original one.
Mistake label classification. We train a multi-modal classifier for mistake label classification. First, the trimmed video clip and step instruction are separately converted into features using EgoVLPv2 [17]. Then, the video features are averaged along the temporal dimension and concatenated to the text features. Finally, the concatenated vectors are forwarded to a two-layer perceptron to predict the logits of the mistake classes. Regarding model training, we compute the class-balanced focal loss [19] because of the relatively low frequency of the mistake and correction classes. We feed the ground-truth step segments to stabilize training whereas the predicted ones are used for testing [20, 21, 22].
4.3 Experiments on mistake action detection
Baseline. We compare our approach with a recent temporal action localization (TAL) model ActionFormer [23] because of TAL’s similarity to our problem settings. [^2] Existing mistake action detection approaches do not fit our settings. For example, the closest one is EgoPED as it predicts both temporal segments and mistake labels [5]. However, its anomaly detection approach predicts binary labels of correct or mistake , thus it cannot solve our three class settings. In our experiments, we train ActionFormer to predict mistake labels for the detected segments, instead of action labels as in TAL.

### v3

@@ Abstract
Figure 1: An example task (cardboard) from our EgoOops dataset. EgoOops includes 50 egocentric videos across five procedural domains and corresponding procedural texts. It contains three types of annotations: video-text alignment, mistake labels, and descriptions explaining the errors in each segment.
@@ 1 Introduction
Table 1: Comparison of mistake datasets. Range: mistake labels to a video or each step segment with start and end times; Cat.: finer-grained categorization than correct , mistake , and correction ; Desc.: descriptions explaining why each segment is incorrect; OM, EM: order mistakes, execution mistakes (see Sec. 3.1); Proc.: workers follow step-by-step procedural texts. Domain specific (e.g., extra screws). Orders unseen in the training set, not always faulty.
@@
However, these studies have the following three limitations (see Tab. 1). L1: video-focused approaches.Existing studies [31, 9, 24, 8, 29, 16, 13] have focused on video-only approaches to mistake detection and not utilized procedural texts. Their approaches are suitable for visually apparent mistakes like incorrectly attached parts and falling plates. Nevertheless, in text-following activities, some mistakes are deviations from procedural texts, thus not obvious only from visual cues (e.g., in Fig. 1, the use of tape is a mistake because the text designates glue). Therefore, besides videos, texts are essential for models to detect mistakes accurately. L2: specific domains.Procedural texts are included in only a few mistake datasets recording cooking [24, 16]. Since many real-world activities follow procedural texts, it is essential to collect data from more domains. L3: rough mistake labels.Most datasets define coarse-grained (correct /mistake /correction[31, 38, 27] or domain-specific (e.g., extra screws[9, 24] categories. General and fine-grained categorization enables analysis of mistake patterns across diverse domains (e.g., a commonly frequent category of mistakes).
@@
We also propose an approach to the problem of mistake action detection, especially focusing on the utilization of procedural texts (L1). To leverage the texts, our approach combines video-text alignment and mistake label classification; the former localizes the start and end times of each procedural step, and the latter assigns the step segments one label of mistake classes. In experiments using EgoOops, our multi-modal approach outperforms a video-only baseline, and the ablation of textual inputs decreases our performance. These results demonstrate that incorporating procedural texts is essential for mistake detection. Additionally, we test existing mistake classifiers and multi-modal large language models for the classification problem. As for the alignment problem, we compare our fully-supervised approach with zero-shot and self-supervised ones. The results confirm that EgoOops targets novel domains of procedural activities and contains useful alignment annotations.
@@ 2 Related work
In this section, we compare EgoOops with other datasets in terms of two perspectives: procedural activity dataset and mistake action dataset.
2.1 Procedural activity datasets
Procedural activity understanding is an important capability for enhancing smart systems, such as AR/VR assistants [38] and intelligent archives. In particular, aligning a sequence of step instructions with a video (i.e., video-text alignment) is fundamental. To support research in this direction, a variety of procedural activity datasets have been developed. While early datasets collected third-person perspective videos accompanied by textual descriptions of each timestep from YouTube [41, 34, 20], recent datasets focus on first-person (egocentric) videos that capture fine-grained details of workers’ activities [33, 17, 5, 22, 26, 10, 1]. For example, EPIC-KITCHENS [5] dataset consists of 432 egocentric videos with step instructions in the cooking domain, including start and end times for each segment. Our dataset provides similar video-text alignment annotations as existing datasets, but distinguishes itself by focusing on erroneous actions, thereby enabling the study of error detection within procedural activities.
2.2 Mistake action datasets
Mistakes are critical in procedural activities, as they can propagate through subsequent steps and significantly affect task success. This has led to the development of various datasets (see Tab. 1). These datasets are categorized into free-style and text-following settings.
@@ 3 EgoOops dataset
EgoOops dataset provides procedural activity videos including erroneous work and annotations for these mistakes. The activities are performed by following instructional steps of procedural texts in order. In this section, we first define mistakes and then describe task selection, video recording, and annotations. Finally, dataset statistics are provided.
@@ 3.1 Mistake definition
We define mistakes as “deviations from instructional steps.” Out of mistake types that meet this definition, we consider executions mistakes in this study. [^1] Other types of mistakes such as order mistakes (e.g., skipping and reordering steps) may also occur during activities, but we focus on the execution mistakes which are our main concern. We leave the study of the other types of mistakes to future work.
Execution mistakes occur when a worker misinterprets and executes steps. We classify the execution mistakes into the six types of errors:
  1. nosep
    Incorrect-object-use (Object ) executes a step with a different object specified in the text. This includes cases when the incorrect number of objects is used.
  2. nosep
    Incorrect-object-picking (Mispick ) picks up an incorrect object, but the worker recognizes the mistake and releases the object. This mistake does not execute a step, unlike incorrect-object-use.
  3. nosep
    Self-correction (Correction ) recognizes and corrects mistakes in a step that has been executed in a wrong way.
  4. nosep
    Accidental-mistakes (Accident ) causes accidental happenings mainly due to carelessness.
  5. nosep
    Wrong-way (Way ) picks a correct object but executes a step in a way that misaligns with the instruction in the text.
  6. nosep
    Other-mistakes (Others ) induces other types of mistakes. This also includes cases where multiple types of the above mistakes occur simultaneously.
3.2 Task selection
To capture a broad range of mistake types described in Sec. 3.1, we construct the EgoOops dataset with activity videos from diverse domains. Accordingly, we select the following five tasks:
  • Electrical circuits (EC ): Connecting electrical elements to complete an electrical circuit that turns a propeller.
  • Color mixture experiments (CM ): Examining the color of various solutions of detergent and fluorescent paint when illuminating them with a blacklight.
  • Ionic reaction experiments (IR ): Examining ionic reaction by dropping chemical solutions to metal plates.
  • Toy building blocks (BB ): Piling up building blocks to construct the specified structure.
  • Cardboard crafts (CB ): Crafting Omikuji boxes, Japanese random fortunes, from cardboard.
We prepare procedural texts for each task as follows: CM and CB use procedures collected from the web; IR and EC use instruction manuals included with out-of-box kits; BB uses a procedure that we manually write from scratch.
@@ 3.3 Video recording
We asked four graduate students to perform the tasks based on the procedural texts. To record their activities, a head-mounted camera (Panasonic HX-A500) was used (Fig. 2). The egocentric videos were recorded at 30 fps with 4K RGB resolution. We chose the egocentric perspective to capture the participants’ visual attention and fine-grained hand-object interactions, which are critical for modeling procedural understanding. During the recordings of EC and BB, images of the final products were given to the participants. [^2] Our preliminary experiments showed these tasks were difficult to complete only with the texts. To avoid the influence of background changes, we fixed the initial locations of the objects, tools, and printed procedural texts. Further, the participants were instructed to work in the chair to capture manipulated objects in detail.
Each participant was asked to perform each task two or four times, totaling 10 recordings per task. For each task, five of 10 recordings contain mistakes that the participants performed intentionally. For the other five recordings, the participants were asked to follow the procedural texts avoiding mistakes. Note that the latter five recordings still contain mistakes due to their careless errors.
Figure 2: Egocentric head-mounted camera on participants.
@@ 3.4 Annotations
EgoOops provides annotations of video-text alignment, mistake labels, and textual descriptions of the mistakes.
Video-text alignment refers to the segments in the recorded videos, and each segment is represented as a start and end timestamp. Each segment corresponds to a human action, which is a step in the procedural text or a mistake (e.g., grasping the wrong object).
Mistake labels are assigned to segments that correspond to execution mistakes. A mistake label is one of the six mistake types in Sec. 3.1. In addition to the mistake labels, we also provide textual descriptions of why the performed actions are considered mistakes.
We asked two persons for annotation, who are called annotator-A and annotator-B to avoid confusion. Annotator-A was asked to annotate the whole dataset using a web annotation tool that we developed. After the annotation, annotator-B was asked to annotate several videos from the dataset again to calculate inter-annotator agreements. For video-text alignment, annotator-B extracted segments from videos and mapped them to the corresponding step labels. The temporal Intersection over the union (tIoU) to the original segments was then calculated. For mistake labels and descriptions, annotator-B was asked to provide mistake labels and descriptions based on the segments annotator-A annotated. The labels and descriptions between the annotators were compared using Cohen’s kappa [3] and BERTScore [40]. The tIOU was 88.888.888.8, and Cohen’s kappa and BERTScore were 86.886.886.8 and 96.396.396.3, respectively, ensuring that the annotations of the dataset are consistent.
@@ 3.5 Statistics
Table 2 shows different trends between the tasks in terms of video duration, segment duration, and the number of segments. For video duration, the longest is the cardboard task at 26.1 minutes, while the shortest is the building block task at 1.9 minutes. As for the texts, Tab. 2 compares the number of steps in a procedural text and the number of words per step. The task with the most steps is the cardboard task, while the task with the longest instructions is the building block task. These statistics indicate that EgoOops covers a variety of procedural tasks ranging from short to long.
@@
#vid.
Avg.
#seg.
#seg.
/ #vid.
Avg.
#steps
per text
#words
/ #steps
@@
Table 3 shows the counts of the labels for execution mistakes. In total, EgoOops contains 95 execution mistakes. Counting the number of each type of mistake, the two most frequent labels are incorrect-object-picking (label 2) and wrong-way (label 5). In addition, we also find unique mistake patterns of each task. For example, accidental-mistakes (label 4) frequently happen in ionic reaction experiments but rarely in the other tasks. We expect the reason is that the ionic reaction experiments involve moving a small metal piece with tweezers and dropping solution into a narrow space. Overall, many execution mistakes occur in the videos of EgoOops, and the tasks have their own frequent mistake types.
Table 3: The number of mistake labels.
@@ 4 Application: mistake action detection
We propose a text-oriented approach to mistake action detection. This section formalizes the problem as consisting of video-text alignment and mistake label classification and explains our approach to them.
@@ 4.1 Problem formalization
Mistake action detection consists of two problems in our formalization: video-text alignment and mistake label classification. It leads to localizing temporal segments and classes of mistakes in a video.
@@
The alignment outputs are passed to the next problem of mistake label classification, which predicts a mistake label for each step segment. The $k$-th step segment is represented by extracted video frames based on the start and end frame numbers as $\mathbf{V}^{\prime}_{k}=(\mathbf{f}_{s_{k}},\dots,\mathbf{f}_{s_{e}})$. Our objective is to assign the segment $\mathbf{V}^{\prime}_{k}$ one label of mistake classes $(c_{1},\dots,c_{N})$, where $c_{n}$ and $N$ represent the name of the $n$-th class and the number of the classes, respectively. In our experiments, we group the mistakes except for 3. correction (see Sec. 3.1) into the common mistake class and solve the classification of three classes: correct , mistake , and correction .  [^3] Our preliminary experiments showed that distinguishing the seven classes (six classes in Sec. 3.1 plus the correct class) is difficult currently.
4.2 Video-text alignment
For video-text alignment, we enhance an existing model of StepFormer [7] by introducing an additional fully supervised loss function, termed StepFormer++. We first provide an overview of the original StepFormer and then explain how we extend it.
Preliminary: StepFormer. StepFormer was originally proposed for learning video-text alignment from untrimmed videos accompanied by narrations in a self-supervised manner [7]. StepFormer is a Transformer decoder [36] equipped with $U$ learnable queries. Given video features extracted using UniVL [19], the queries attends to the video features, producing $U$ contextualized vectors called step slots , which capture key steps in the video. The step slots temporally align with narration vectors extracted using UniVL, where they use a sequence-to-sequence alignment algorithm Drop-DTW [6]. Considering this alignment as positive pairs, the loss to supervise the step slots is calculated as contrastive one InfoNCE [35] at both local (same video-narration pairs) and global (different video-narration pairs) levels. During inference, StepFormer can temporally localize the video segment of each step instruction (i.e., video-text alignment). Specifically, the extracted step slots align with the step instructions in procedural texts, allowing unmatched slots to be dropped. Next, the remaining slots are aligned with the video to identify the start and end times of each step. In these alignment processes, Drop-DTW is used again.
Pre-training. We select StepFormer because its self-supervised learning can be used for pre-training to mitigate the negative impact of the small size of EgoOops. In the original paper [7], StepFormer is trained on untrimmed web videos and narrations in HowTo100M [20] using UniVL [19] features. Instead, we train it on Ego4D [10] using EgoVLPv2 [25] features to fill the domain gap between web and egocentric videos. Ego4D is a massive-scale egocentric video dataset accompanied by transcriptions [10], hence we can pre-train StepFormer following the same procedure as the original one.
StepFormer++. The pre-trained model is fine-tuned on EgoOops with an additional loss to train StepFormer in a fully supervised setting by leveraging the video-text alignment annotations. Figure 3 shows an overview of the resulting model StepFormer++. The overall process is the same as the original StepFormer. Given $(\mathbf{V},\mathbf{T})$, the model first extracts video $\mathbf{H}_{v}=(\mathbf{h}_{v}^{1},\dots,\mathbf{h}_{v}^{l},\dots,\mathbf{h}_{v}^{L})$ and text $\mathbf{H}_{t}=(\mathbf{h}_{t}^{1},\dots,\mathbf{h}_{t}^{k},\dots,\mathbf{h}_{t}^{K})$ features using EgoVLPv2 instead of UniVL. Then, the Transformer decoder makes slot queries attend $\mathbf{H}_{v}$, producing step slots $\mathbf{S}=(\mathbf{s}_{1},\dots,\mathbf{s}_{u},\dots,\mathbf{s}_{U})$. Finally, the model acquires the step-text alignment via Drop-DTW by computing a similarity matrix of $\mathbf{S}$ and $\mathbf{H}_{t}$. The step-text alignment is used to compute the original loss at only the global (different video-narration pairs) level. Instead of the local (same video-narration pairs) loss, we add a new loss to the StepFormer for learning from video-text alignment annotations. After Drop-DTW between the slots and step instructions [^4] We leave the same number of slots as the steps in the procedural text., we supervise the remaining step slots $\hat{\mathbf{S}}=(\hat{\mathbf{s}}_{1},\dots,\hat{\mathbf{s}}_{k},\dots,\hat{\mathbf{s}}_{K})$ to match with the start and end frame numbers $(s_{k},e_{k})$. Specifically, this is calculated using the InfoNCE framework:
$\mathcal{L}_{\mathrm{supervised}}(\hat{\mathbf{s}}_{k},\mathbf{H}_{v})=-\log\frac{\sum_{j\in[s_{k},e_{k}]}f(\hat{\mathbf{s}}_{k},\mathbf{h}^{j}_{v})}{\sum_{l}f(\hat{\mathbf{s}}_{k},\mathbf{h}^{l}_{v})},$ (1)
where $f(\hat{\mathbf{s}}_{k},\mathbf{h}^{*}_{v})\!=\!\exp(\cos(\hat{\mathbf{s}}_{k},\mathbf{h}^{*}_{v}))/\gamma$, and $\gamma$ is a scaling temperature. We add this loss to the original ones, and the pre-trained model is fine-tuned on EgoOops.
Figure 3: An overview of StepFormer++.
4.3 Mistake label classification
Multi-modal classifier. Given a pair of predicted video segment $\mathbf{V}^{\prime}$ and $k$-th step instruction $\mathbf{t}_{k}$, the model predicts the mistake label. Specifically, the model first convert $\mathbf{V}^{\prime}$ and $\mathbf{t}_{k}$ into video $\mathbf{H}_{v}^{\prime}=(\mathbf{h}_{v}^{s_{t}},\dots,\mathbf{h}_{v}^{e_{t}})$ and text $\mathbf{h}_{t}^{k}$ features using EgoVLPv2. Then, it computes the mean of $\mathbf{H}_{v}^{\prime}$, concatenates the averaged vector with $\mathbf{h}_{t}^{k}$, and forwards it into a two-layer perceptron $g$ with ReLU function as following: $\mathbf{z}_{k}=g(\mathrm{concat}(\mathrm{mean}(\mathbf{H}_{v}^{\prime}),\mathbf{h}_{t}^{k}))$, where $\mathbf{z}_{k}=(z_{k}^{1},\dots,z_{k}^{n},\dots,z_{k}^{N})$ represents the logits for classes, and $N$ is the number of the classes. The model applies the argmax operation on $\mathbf{z}_{k}$ and outputs the prediction label.
Training. To train the model, we use the class-balanced focal loss [4] because the frequency of the mistake and correction labels is lower than the correct label. Specifically, using $\mathbf{z}_{t}$, the loss is calculated as following:
$\mathcal{L}_{\mathrm{classification}}(\mathbf{z}_{t})=-\frac{1-\beta}{1-\beta^{r_{c_{n}}}}\log\frac{\exp(z_{t}^{c_{n}})}{\sum_{j}\exp(z_{t}^{c_{j}})},$ (2)
where $r_{c_{n}}$ is the number of training samples belonging to the class $c_{n}$, and $\beta\in[0,1)$ is a hyperparameter. We adopt teacher forcing [11, 36] as the training strategy. Specifically, we input the ground-truth segment of the $t$-th step to the model to stabilize training whereas the predicted ones are used for testing.
4.4 Implementation details.
We follow the official implementation of StepFormer and use the same hyperparameters as stated in [7] unless we mention modifications. We set the number of step queries to be $U=32$ and the batch size to be 6 for fine-tuning StepFormer++ on EgoOops. The video and text feature dimension of EgoVLPv2 is $d=4,096$. We use Drop-DTW with an 80 percentile drop cost [6] to align the step slots and video features. We set $\gamma=0.03$ in the InfoNCE loss and $\beta=0.9999$ in the class-balanced loss. For the mistake label classification, we train the classifier in 1,200 epochs.
5 Experiments
We first report an end-to-end performance on mistake action detection in Section 5.1. We then conduct in-depth experiments on video-text alignment and mistake label classification individually in Section 5.2 and 5.3.
5.1 Mistake action detection
Baseline. We do not adopt existing mistake detection methods as baselines because they do not fit our settings. For instance, Assembly101 [31] and CaptainCook4D [24] assume trimmed video clips as inputs, while our task assumes untrimmed videos. The method in [29] focuses on ordering mistakes in the videos and does not address execution mistakes. EgoPED [16] and AMNAR [13] are the closest to our setting as they predict both segments and mistake labels. However, their methods are based on anomaly detection, predicting binary labels of “correct” and “mistakes,”; thus they cannot predict the three classes of “correct,” “mistakes,” and “correction.”
Therefore, instead of existing mistake action detection models, we compare our method with the recent temporal action localization (TAL) model, ActionFormer [39]. This is because TAL operates under similar conditions, where the models detect both temporal segments and their action labels. In our experiments, we train ActionFormer to predict mistake labels for the detected segments, instead of action labels as in the original settings. For a fair comparison, we apply NMS to retain as many segments as the ground truths. In addition, since our metrics require step labels, we assign them to the segments in order from the start to the end of the videos. Note that the inputs for TAL are videos only.